A moving coil instrument gives a full scale deflection (F.S.D) with a current of 30mA (0.030A) and a voltage of 90mv(0.090v). Calculate the value of a resistor to be connected in series with the instruments so that it can be used to read 0-100voltsr

2)__ A moving coil instruments gives a full scale deflection with a potential difference of 70mV and a current of 20mA calculate the value of the shunt required to give a range of 0-50Amps.
1
Expert's answer
2020-02-06T08:24:47-0500

1)


"Vo=V-Vi=100V-0.09V=99.91V"

"Ro=Vo/Ii=99.91V/0.03A=3330.(3) \Omega"

Answer: 3330om

2)


"Io=I-Ii=50A-0.02A=49.98A"

"Ro=Vi/Io=0.07V/49.98A=0.0014\Omega"

Answer: 0.0014om